Status Report

NASA Space Station On-Orbit Status 12 February 2005

By SpaceRef Editor
February 14, 2005
Filed under , , ,
NASA Space Station On-Orbit Status 12 February 2005

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2005) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below. Saturday, first weekend rest day for the crew.

CDR/SO Leroy Chiao and FE Salizhan Sharipov began the day with the regular weekly 3-hr. task of thorough station cleaning. [“Uborka”, done every Saturday, includes removal of food waste products, cleaning of compartments with vacuum cleaner, wet cleaning of the Service Module (SM) dining table and other surfaces with “Fungistat” disinfectant and cleaning fan screens to avoid temperature rises.]

Chiao disconnected and removed the UOP DCP (utility outlet panel/display & control panel) bypass power cable at the Lab Robotics Work Station, which was used to support video camera coverage of yesterday’s SSRMS (Space Station Remote Manipulator System proficiency training ops.

Leroy was congratulated on his great job on the SSRMS training, which was a complete success. [There were no issues with the new MSS (Mobile Service System) R3.1 software, and the grapple fixture did not stick. The SSRM was placed back into EVA-12 viewing position. The ground got useful data for the “sticky grapple fixture release” phenomenon and continues to process all the data to further investigate the options for overcoming the problem, if it should crop up again.]

The CDR took did his daily readings of Total Dose and End File values of the TEPC (Tissue Equivalent Proportional Counter), which he had relocated to the Node on 1/31, and called the data down at the evening DPC (daily planning conference). [This is currently a daily requirement since the UOP (utility outlet panel) near the TEPCs temporary location (for two weeks) does not allow automated telemetry monitoring by the ground.]

In the Service Module (SM), the FE completed the weekly SOZh/ECLSS (environment control & life support systems) servicing tasks, which today were extended for some non-routine IFM (inflight maintenance) activities. [For some time overnight and today, ground specialists at TsUP/Moscow and station crew (wearing face masks, goggles and gloves) performed troubleshooting on the ASU toilet facilities, which had a backflow and contamination problem with the preservative/pre-treat solution in the hoses up to the waste separator and fan. For the duration of the troubleshooting, the crew was cleared to use the Soyuz TMA-5 waste collection system (which has a one-day reserve capability). The E-K preservative container normally holds five liters of Tox Level 2 pre-treat solution, i.e., a mix of H2SO4 (sulfuric acid), CrO3 (chromium oxide, for oxidation and purple color), and H2O (water). The pre-treat liquid is mixed with water in a dispenser (DKiV) and used for toilet flushing.]

The crew worked out in accordance with their daily 2.5-hr. physical exercise protocol on TVIS treadmill (aerobic), RED exerciser (anaerobic) and VELO cycle with bungee cord load trainer (aerobic plus anaerobic). [Salizhan’s daily protocol currently prescribes a four-day microcycle exercise with 1.5 hr on the TVIS (today: Day 1 of a new set) and one hour on VELO.]

Leroy then transferred the daily TVIS and RED exercise data files to the MEC (medical equipment computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data, followed by their erasure on the HRM storage medium.

For his “Saturday Science” program today, the Science Officer worked on the HRF FOOT (Foot/Ground Reaction Forces during Space Flight) experiment, transferring, consolidating and configuring the FOOT Hardware Supply Kit, including the specially instrumented LEMS (lower extremity monitoring suit) pants garment. [This provides a universal kit for coming FOOT operations, thereby reducing future upmass requirements and allowing for excess hardware to be removed from Station.]

Working off the discretionary Russian task list, Sharipov downloaded Molniya-SM/LSO experiment measurements recorded during the last session (9/26-28, 2004) from the EGE-1/LSO memory to HDD (hard-disk drive) on the EGE-2 laptop for return to Earth. He then disassembled the Molniya hardware at SM window #3 and stowed it. [Objective of Molniya-SM, similar to the French LSO experiment, was to catch and record incidental storm phenomena and other related events in the Earth’s equatorial regions. The experiment is controlled from the EGE-1 laptop, loaded with orbital sighting predictions using an up-to-date NORAD tracking TLE (two-line element) provided by NASA. Objective of LSO was to study rare optical phenomena occurring in the upper layers of Earth’s atmosphere, so-called “sprites” (i.e., puzzling glow phenomena observed above thunderstorm clouds). LSO was originally part of Claudie Haigneré’s French “Andromeda” payload package of taxi mission 3S that could not be performed as planned during Increment 4 due to an ISS flight attitude conflict, and also part of André Kuipers’ VC-6 “Delta” science program.]

Leroy Chiao completed the regular weekly maintenance reboot on the operational PCS (portable computer system) laptops and the restart of the OCA comm router laptop (every two weeks).

The crew was given the Go to use the PUL (Portable Utility Light), which was under investigation after the trip of RPC-4 (Remote Power Controller #4) in a Lab RPCM (RPC Module) on 11/11/04. Since then, UOP-4 (Utility Outlet Panel 4) and its loads, which included the PUL, have been cleared one by one.

The crew was thanked for yesterday’s highly successful ADUM (Advanced Diagnostic Ultrasound in Micro-G) scanning. [All objectives were accomplished, and some secondary objectives were obtained as well. The Expedition 9 crew attended the session on the ground.]

At ~9:00am EST, Chiao and Sharipov engaged in their weekly planning conference (WPC) with the ground, discussing next week’s “Look-Ahead Plan” (prepared jointly by MCC-H and TsUP/Moscow planners), via S-band/audio, reviewing upcoming activities and any concerns about future on-orbit events.

Later, at ~9:40am, the crew held their weekly teleconference with ISS Program Management at JSC/Houston via S-band/audio.

Weekly Science Update (Expedition Ten — 15th):

GASMAP: Nothing new.

Human Research Facility/Workstation (HRF WS): Continuing.

Advanced Ultrasound (ADUM): The ADUM team thanked the crew for their hard work this week. “It was quite an adventure with all of the hiccups with the PC, but your perseverance really paid off!” The ground captured great images and is looking forward to discussing them with the crew you next week.

Hand Posture Analyzer (HPA: Nothing new.

In-Space Soldering Investigation (ISSI): Operations are complete.

Pore Formation and Mobility Investigation (PFMI): Nothing new.

Space Acceleration Measurement System (SAMS): Nothing new.

Microgravity Acceleration Measurement System (MAMS): Nothing new.

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES): PCG-STES is performing nominally.

Protein Crystal Growth Monitoring by Digital Holographic Microscope (PromISS): Nothing new.

Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE): Nothing new.

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

Binary Colloidal Alloy Test-3 (BCAT-3): The Science Officer was thanked for his “very clear” pictures of the BCAT-3 samples. Over 240 MB of zipped (compressed) files (50 DCR photos and 17 WAV sound files) were downlinked and are being looked at and analyzed. The SO could not see any surface crystals yet in samples 8, 9, and 10 when he was taking the photographs, so investigators suspect that it’s too soon for any surface crystals to have formed since homogenization. A few weeks ago they transferred the BCAT-3 ground-based sample 10 into a square sample cell, as the one in BCAT-3). It had crystallized on the ground in its cylindrical sample cell over a period of many months, but the transferred sample has yet to crystallize again. The original CDOT glovebox experiments showed that colloidal crystals form in weeks in space when the same sample do not crystallize over a period of years on the ground. This is because the particles do not jam since there is a lack of sedimentation (caused by gravity) on the ISS.

Renal Stone (RS): Nothing new.

Synchronized Position Hold, Engage, Reorient, Experimental Satellites (SHERES): Nothing new.

Foot/Ground Reaction Forces during Space Flight (FOOT): The Science Officer was thanked for taking time out of his Saturday to configure the Foot hardware kit. “This will go a long way towards ensuring Foot’s success in future increments”.

Materials ISS Experiment (MISSE): In progress. Deployed outside on the U.S. Airlock. Nominal and collecting data.

Cellular Biotechnology Support Systems-Fluid Dynamics Investigation (CBOSS-FDI): Nothing new.

Dust and Aerosol Measurement Feasibility Test (DAFT): Nothing new.

Biotechnology Specimen Temperature Controller (BSTC): Nothing new.

Yeast Group Activation Packs (Yeast GAP): Nothing new.

Earth Knowledge Acquired by Middle School Students (EarthKAM): Nothing new.

Earth Science Toward Exploration Research (ESTER): Nothing new.

Serial Network Flow Monitor (SNFM): Nothing new.

Fluid Merging Viscosity Measurement (FMVM): Looking forward to future operations with honey samples.

Space Experiment Module (SEM): The SEM team thanked the crew again with a message from the students at Mott Hall School in N. Harlem, NY. The students are eager to share their story and express their gratitude for your hard work.

Viscous Liquid Foam–Bulk Metallic Glass (Foam): Nothing new.

Effects of Prolonged Space Flight on Human Skeletal Muscle (BIOPSY): Nothing new.

Coarsening in Solid-Liquid Mixtures-2 (CSLM-2): Planned.

Commercial Generic Bioprocessing Apparatus (CGBA): Nothing new.

Miscible Fluids in Microgravity (MFMG): Nothing new.

Educational Payload Operations (EPO): Nothing new.

Capillary Flow Experiment (CFE): Nothing new.

Crew Earth Observations (CEO): The ground has to date received a total of over 10,000 CEO images during Increment 10, and the weekly volume continues to increase. Investigators have received and preliminarily reviewed imagery through 2/9. In general, the crew’s long-lens views are of excellent quality. “Your beautiful shots of Singapore are the most detailed we have ever seen. Nice work!”

Today’s CEO (crew earth observations) photo targets, limited in the current XPOP attitude by flight rule constraints on the use of the science window, which is available for only ~1/4 of each orbit when not facing forward (in ram), were Mekong River Delta (this overpass provided an opportunity for sunglint illumination of river channels in the Mekong Delta. If sunglint was visible [to the right of track], it greatly improved the ability to recognize small channels and embayments. Mapping of these small water courses is useful to understand changes in estuarine dynamics resulting from land use changes in the delta region), and Internal Waves, North Atlantic (a clear weather pattern is holding to the west of the Iberian Peninsula. This overpass provides an opportunity to photograph internal waves to the north of the Azores. Looking to the right of track for the sunglint point).

CEO photography can be viewed and studied at the websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 10 crew visit:

Expedition 10 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

Upcoming Key Events:

  • ISS Reboost — 2/15 (~8:22pm EST, delta-V ~1.8 m/s; delta-H ~3.2 km; phasing for 17P launch);
  • Progress 16 (16P) undocking & destructive reentry — 2/27;
  • Progress 17 (17P) launch — 2/28;
  • EVA-13 — 3/25;
  • Soyuz 10 (10S) launch — 4/15 with Expedition 11 (CDR Sergei Krikalev, FE/SO John Phillips);
  • Soyuz 9 (9S) undock — 4/25 with Exp. 10 crew (after 193 days on orbit, 191 days on board ISS);
  • Progress 18 (18P) launch — 6/10;
  • Progress 19 (19P) launch — 8/24;
  • Soyuz 11 (11S) launch — 9/27.

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

ISS Orbit (as of this morning, 6:18am EST [= epoch]):

  • Mean altitude — 356.4 km
  • Apogee height — 363.0 km
  • Perigee height — 349.9 km
  • Period — 91.67 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.000974
  • Solar Beta Angle — -33.6 deg (magnitude decreasing)
  • Orbits per 24-hr. day — 15.70
  • Mean altitude loss in last 24 hours — 130 m
  • Revolutions since FGB/Zarya launch (Nov. 98) — 35607

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.